Three times the larger of the two numbers is 10 more than twice the smaller. Five times the smaller is 11 less than four times the larger. What are the numbers?
I need help setting up the two equations.

Answer by ankor@dixie-net.com(22740)

let x = the larger number
let y = the smaller number
:
Write an equation for each statement:
:
"Three times the larger of the two numbers is 10 more than twice the smaller."
3x = 2y + 10
:
" Five times the smaller is 11 less than four times the larger."
5y = 4x - 11
:
Put both equation in the same form,
to use elimination, Mult the 1st by 5, the 2nd by 2 and add
3x - 2y = 10
-4x + 5y = -11